A robot lands on Mars. Its x and y coordinates are given by equations: x(t) = 2m –(0.25m/s2 )t2

y(t)= (1.0 m/s)t +(0.025m/s3)

(a).What is the position and distance from landing at t=2 seconds?(b)What is the displacement and average velocity during the interval from t=0 s to t= 2?(c).Derive a general expression for its instantaneous velocity vector and find its instantaneous velocity in its component form as well as in terms of its magnitude.


1
Expert's answer
2021-10-10T16:04:57-0400

(a) The position and distance are


"x(t) = 2 –(0.25)t^2,\\
x(2)=1\text{ m},\\
y(t)= (1.0)t +(0.025)t^3,\\
y(2)=2.2\text{ m},\\
d=\sqrt{y^2+x^2}=2.42\text{ m}."

(b) The displacement is d, or 2.42 m, the average velocity is


"v=\frac dt=1.21\text{ m/s}."

(c) In component form:


"v_x(t)=x'(t)=-0.5t,\\
v_y(t)=y'(t)=1+0.075t^2.\\
|v|=\sqrt{v_x^2(t)+v_y^2(t)}=\\
=\sqrt{1+0.4t^2+5.625·10^{-3}t^4}."